Apr 6 (UT)

Observing log goes here

Observers: Christina Williams, Zheng Cai, Ian McGreer

2:00 UT Dome opened (for example).

01:03 UT Taking darks

02:26 UT Moving to first telluric (williams)

02:28 UT Telescope pointing is off

02:44 UT Pointing back on, trying again. Didn't work

02:57 UT Not fixed. Trying target field instead (williams)

03:22 UT Fixed pointing on target field (williams). Finishing up the fix.

03:28 UT Trying 1st telluric again (williams) failed. Unclear why. Moving to science field cause the guide star worked there.

03:37 UT Starting MOS acquision (williams)

04:06 UT Starting Science script (williams) Avg seeing = 0.76"

04:39 UT Clouds. Starting with file 30,31. 3 mag of extinction.

04:50 UT Guiding is lost. Starting with file 32.

05:11 UT Extinction went down for files 34,35ish

05:14 UT starting 2nd run science script (williams)

05:24 UT file 38,39 affected by clouds

05:40 UT file 41,42,43 affected by high extinction

06:02 UT starting with file 44 looks better.

07:28 UT finished 3rd script, moving to telluric. Wind speeds high

07:30 Pointing problem again. Doing a pointing check nearby.

07:58 UT 2nd Telluric running (williams)

08:05 UT Going to imaging target, avg seeing 0.65 (McGreer) Pointing is off.

08:09 UT Presets (McGreer)

08:11 UT Science script running (McGreer) seeing = 0.5-0.6

08:26 UT High extinction, patches of clouds

08:29 UT Lost guiding. Resent aquisition

08:53 UT DIM is up and running. says seeing is 1.1. Just kidding, it changed to 0.8. Guider says seeing is 0.5-0.7

09:13 UT McGreer imaging completed. Guider seeing 0.7

09:16 UT Starting Cai acquisition script. Pointing problem again (no guidestar found)

09:18 UT Running pointing check

09:22 UT Re-sending Cai acquisition script

09:41 UT Running Cai acquisition. seeing =0.9", a little cloudy

09:45 UT Lost guiding. Continuing with aquisition

09:46 UT Guiding is back.

10:05 UT Starting spectroscopy script (Cai)

10:22 UT Guiding temporarily lost, clouds, file 99 is affected

10:30 UT File 101 affected by extinction.

11:14 UT Running another 1/3 of Cai script

12:04 UT Stopping Cai science, moving to Cai telluric

12:16 UT Starting telluric acquisition (Cai)

12:24 UT Starting telluric spectrum (Cai)

12:34 UT Done with Cai telluric. Getting ready for Cals

12:45 UT Running Flat

13:17 UT Calibrations complete

-- %USERSIG{log author - 2015-04-06}%

Comments


 
Topic revision: r4 - 06 Apr 2015, EiichiEgami
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback